# SegVol: Universal and Interactive Volumetric Medical Image Segmentation

### v3

@@ 2.2 Internal validation compared with task-specific segmentation models
Figure 3: Quantitative results of external validation experiments. a. Violin plots for comparison experiment results of SegVol and interactive methods. The vertical axis represents the Dice score. b. The bar chart illustrates the consistency of SegVol’s performance on the external validation set across different prompt types.
2.3 External validation compared with interactive methods
@@
Several efforts have been made to construct an interactive segmentation model. However, some of these works, such as MedSAM[38] and SAM-MED2D[6], focus on 2D tasks and cannot process 3D input directly. The other 3D-based methods, such as SAM-MED3D[64], only support small cropped input and lack semantic information support, which is still far from building a comprehensive foundation model for volumetric medical image analysis. To compare with these interactive segmentation models, we performed external validation experiments on 1,738 cases from the validation set of AMOS22[26] and the whole novel annotated set of Universal Lesion Segmentation Challenge 23(ULS23)[9]. The validation set of AMOS22 contains 120 cases annotated with 15 main organs. The novel annotated ULS23 dataset is composed of three subsets, namely, DeepLesion3D, Radboudumc Bone, and Radboudumc Pancreas. The DeepLesion3D subset contains 200 abdominal lesions, 100 bone lesions, 50 kidney lesions, 50 liver lesions, 100 lung lesions, 100 mediastinal lesions, and 150 assorted lesions cases. There are 744 bone lesion cases in the Radboudumc Bone subset and 124 pancreas lesion cases in the Radboudumc Pancreas subset.
The quantitative results of external validation experiments are shown in Fig. 3. The Fig. 3 a illustrates our method is the best in most of the tasks including lesions and organs, compared to other SAM-like interactive models. MedSAM[38] and SAM(bounding box)[31] use bounding box prompts. SAM(5 clicks)[31], SAM-MED2D[6] and SAM-MED3D[64] use point prompts and a five-step correction procedure, which means that the point prompt in each step will be given according to the previous-step output and ground truth, rather than giving all at once. In this experiment, our SegVol uses bounding box and text prompt which performs better than other kinds of prompt combinations. We provide the ablation study in Fig. 3 b, which shows the good performance of SegVol among different prompt types, especially bbox(bounding box) prompt and text+bbox prompt. Note that the category of each mask in ULS23 is not clearly defined. Thus, we give a general text (i.e. ‘tumor’ or ‘lesion’) to prompt SegVol, and it is compatible with such general prompts and performs well. The detailed scores and visualization results of external validation are presented in Supplementary Table 5 and Fig. 6-8.
In Fig. 4, we visualize the segmentation results in 4 important organ categories to study the differences within these interactive models. Due to the lack of understanding of the 3D structure, 2D methods like MedSAM and SAM(bbox) present worse results. Although the 3D segmentation method, SAM-MED3D, performs well in the easy aorta case, it demonstrates poor segmentation results in others, especially the pancreas and stomach cases. SegVol achieves good segmentation results stably in all categories, relying on its full understanding of 3D spatial structures and semantics.
In addition, we discuss the generalization performance of SegVol on an external MRI dataset. We collect 60 MRI scans annotated with 4 key organ categories from CHAOS[28, 30, 29] dataset and evaluate the generalization ability to unseen modality of SegVol. It achieves median Dice scores of 85.70%, 80.09%, 80.04%, and 81.46% for liver, spleen, left kidney, and right kidney, respectively. This generalization result demonstrates the robustness of SegVol in the face of completely unseen modality data. The detailed scores and visualization results are presented in Supplementary Table 4 and Fig. 9.
Figure 4: Visualized prediction results of SegVol and other interactive methods on four external categories. In each case, the upper row is the axial plane of CTs and the lower row is the sagittal plane of CTs. Visualization results show the accuracy and robustness of SegVol’s modeling of various anatomical structures.
Figure 5: Analysis of the relationship between semantic-prompt and spatial-prompt. a. The quantitative experimental results on 19 internal tasks demonstrate that jointly using semantic and spatial prompts can achieve better performances. b. The four cases demonstrate that semantic prompts can clarify the ambiguity of spatial prompts and avoid multi-plausible outputs. Each image shows the segmentation result of SegVol using the spatial prompt, i.e. point or bounding box, and semantic prompt, i.e. the caption below the image. c. We reflect the spatial prompts to semantic categories (prompts). Each image shows the spatial prompt and the mask prediction. The bar charts rank the top 8 semantic categories with the highest probabilities.
2.4 The interaction relationship between spatial-prompt and semantic-prompt
As a universal model, our approach achieves precise segmentation for over 200 organs, tissues, and lesions using both spatial and semantic prompts. In Fig. 5 a, we quantitatively analyze the mutually supportive relationship between semantic-prompt and spatial-prompt in 19 internal segmentation tasks. On the one hand, spatial prompts allow the model to locate the specific part in the 3D space. According to Fig. 5 a, the average Dice score of ‘bbox+text’ prompt is boosted by 5.85% compared to the ‘text’ prompt on average. On the other hand, semantic prompts clarify the reference to the anatomical structure, eliminating the ambiguity of spatial prompts and the plausible masks of multiple categories. This is reflected in Fig. 5 a as the average Dice score of ‘point+text’ prompts is 4.62% higher than using ‘point’ prompts alone. Spatial and semantic prompts mutually support each other, ultimately endowing the model with powerful segmentation capabilities.
Alexander Kirillov, etc.[31] discuss the multiple plausible outputs problem in the spatial prompt setting. As illustrated in the images on the top left in Fig. 5 b, they are applied with the same point prompt which reasonably corresponds to three concepts, namely, kidney tumor, left kidney, and the whole kidney. Similarly, in the bottom left images, the bounding box selects the region of liver. However, liver tumors, hepatic vessels, and liver itself are also plausible target structures. In these cases, SAM chooses to return multiple masks to match different levels of plausible results. Unlike SAM’s solution, we use semantic prompts to clarify the targets. As shown in Fig. 5 b, the captions below the images are text prompts, and the masks in the images are the predictions of SegVol, which shows that the semantic prompts can effectively disambiguate the text prompts.
Furthermore, we study the possibility of SegVol to reflect spatial prompts to semantic categories. Fig. 5 c reveals that SegVol can give accurate semantic categories based on the spatial prompts. In the top left image in Fig. 5 c, the spatial prompt on the liver results in a 0.997 prediction score for liver. The top right image in the sub-figure shows if the spatial prompt is the point on the liver tumor, SegVol will output a 0.619 prediction score for tumor category and a 0.339 prediction score for liver based on the spatial relationship of liver tumor and liver. We implement this reflection experiment by decoding the semantic prompts from a category set and applying the softmax function among the logits of semantic prompts on the predicted mask voxels to get the prediction probabilities of different categories.
2.5 Scaling up training data
@@
The success of scaling up has been witnessed in multiple computer vision tasks [31, 45]. We conduct an ablation study to investigate the importance of scaling up training images and masks. The BTCV dataset[32], which includes 13 main organs, is set as an anchor to evaluate the model trained separately on 1, 2, and 8 datasets for 500 epochs, as well as the final model trained on 25 datasets. The detailed results are shown in Fig. 2 b. As a lightweight model, the performance is weak when only one dataset is used. However, with the increase of training data, the Dice score increases rapidly, especially in the text prompt setting. The results indicate that our method is scalable and better performance can be achieved if more training data is available.
@@ 3 Discussion
We present SegVol, a 3D foundational model for interactive and universal volumetric medical image segmentation. This method has been developed and evaluated using 25 open-source datasets, 10 internal validation tasks, and 18 external validation tasks. Unlike the traditional volumetric segmentation method, nnU-Net[24], which automatically configures settings for every dataset, SegVol is designed to unify various volumetric segmentation datasets into a single architecture. This results in a universal segmentation tool capable of generating accurate responses for over 200 anatomical targets. Furthermore, SegVol demonstrates state-of-the-art volumetric segmentation performance when compared with both traditional task-specific methods[15, 61, 24, 33] and the recent interactive methods[38, 6, 64, 31] in internal validation and external validation experiments, respectively. Despite its universality and high precision, SegVol maintains a lightweight architecture compared to other volumetric segmentation methods. We have made SegVol an open-source foundational model, readily applicable to a broad spectrum of medical image representation and analysis fields. This ensures it can be easily integrated and utilized by researchers and practitioners alike.
While SegVol is capable of understanding semantic prompts composed of sentences, there remains a gap between it and the referring expression segmentation that involves complex semantic information and logical relationships. The establishment of a referring expression segmentation model needs more curated data with spatial annotations with text. Our SegVol provides a foundation for realizing referring segmentation of medical images, and we leave as future work.
We primarily use SegVol with Computed Tomography (CT) data due to its advantages of easy acquisition, wide usage, and high resolution. CT is also the preferred method for evaluating solid tumors. Furthermore, the flexible architecture of SegVol allows it to be compatible with various types of volumetric medical images, like MRI. The current framework also allows for the direct addition of new training data in the same format, even if the new data are from unseen categories. This means that the model can inherit all previous knowledge and continue learning in new fields. This adaptability and continuous learning capabilities make SegVol a promising and broadly used tool in the field of medical image analysis.
4 Method
4.1 Data processing
One of the main challenges in volumetric medical image segmentation is the absence of large-scale publicly available volumetric medical data, especially the annotated segmentation CTs. Doing our utmost, we collected 25 open-source segmentation CT datasets, including CHAOS[28, 30, 29], HaN-Seg[44], AMOS22[26], AbdomenCT-1k[39], KiTS23[22], KiPA22[19, 18, 55, 56], KiTS19[21], BTCV[32], Pancreas-CT[49, 50, 8], 3D-IRCADB[59], FLARE22[58, 40], TotalSegmentator[65], CT-ORG[47, 48, 2, 8], VerSe19, VerSe20[53, 36, 34], SLIVER07[20], QUBIQ[1], six MSD datasets[58], LUNA16[54], and WORD[37]. These CTs originate from various medical institutions, captured by different machines with varying parameter settings and scanning regions. These factors result in a wide data distribution, and thus significant challenges in data processing.
To standardize these datasets, we perform the following transformation on every CT scan. Firstly, we set a threshold based on the mean voxel value of each volume. Voxels with values that are above this threshold are retained. Then, we calculate the $99.95^{\text{th}}$ and $0.05^{\text{th}}$ percentiles of the remaining voxels and use them as the upper and lower bounds to clip the original voxels and obtain the foreground. Finally, we normalize the foreground voxels using the mean and standard deviation.
Volumetric segmentation datasets suffer from the notorious problem of partial labels. Most of the datasets have annotations of only a few segmentation targets, e.g., several organs. Therefore, the deep models may learn the spurious correlation between datasets and segmentation targets, and produce inferior results during the inference phase. To relieve this problem, we introduce the pseudo labels by utilizing the Felzenswalb-Huttenlocher (FH)[11] algorithm to generate pseudo masks for each CT scan.
The unsupervised segmentation algorithm FH[11] separates the spatial structures based on the gradient between adjacent voxels. However, pseudo masks derived by the FH algorithm contain substantial noise and numerous small masks, for example, the disconnection of a complete structure and the wrong connection of different structures. To improve the pseudo masks, we employ the following strategies: 1) The pseudo masks are replaced with ground truth masks when applicable. 2) We filter out tiny structures smaller than 1‰ of the whole volume. 3) Each mask is refined by dilation and erosion operations.
4.2 Model architecture
The volumetric medical image segmentation dataset $\mathcal{D}=\{(\bm{x}_{i},\bm{y}_{i})\}$ consists of many pairs of 3D images and mask labels. Each data pair has a 3D image datum $\bm{x}_{i}\in\mathcal{R}^{C\times D\times H\times W}$ and $K$ mask labels $\bm{y}_{i}\in\{0,1\}^{K\times D\times H\times W}$, corresponding to $K$ target categories. The classic segmentation model[16, 71, 15, 61, 24, 33] $\mathcal{F}(*,\bm{\theta})$ learns to predict masks $\bm{y}_{i}$ belonging to the $K$ categories based on the volumetric input $\bm{x}_{i}$, i.e., $\bm{o}_{i}=\mathcal{F}(\bm{x}_{i},\bm{\theta})$, where $\bm{o}_{i}\in\mathcal{R}^{K\times D\times H\times W}$. Therefore, the traditional models are not able to generalize to unseen categories.
Motivated by the recent advance in 2D nature image segmentation, Segment Anything (SAM)[31], we design a novel method for interactive and universal volumetric medical image segmentation, named, SegVol. We illustrate the model in Fig. 1 a. SegVol supports three types of prompts for interactive segmentation: ‘bbox’ prompt, $\bm{b}\in\mathcal{R}^{6}$ representing the coordinates of two diagonal vertices; ‘point’ prompt, including a set of ($P$) points $\bm{p}\in\mathcal{R}^{P\times 3}$; and ‘text’ prompt, such as ‘liver’ or ‘cervical spine C2’, which is tokenized to tensor $\bm{t}$. SegVol consists of four modules, namely, image encoder $\mathcal{F}_{\text{IE}}(*,\bm{\theta}_{\text{IE}})$, text encoder $\mathcal{F}_{\text{TE}}(*,\bm{\theta}_{\text{TE}})$, prompt encoder $\mathcal{F}_{\text{PE}}(*,\bm{\theta}_{\text{PE}})$, and mask decoder $\mathcal{F}_{\text{MD}}(*,\bm{\theta}_{\text{MD}})$. We introduce each module in the following.
We employ ViT (Vision Transformer)[10] as the image encoder, which exhibits remarkable advantages over convolutional models[17] when pre-trained on large-scale datasets. We first pre-train ViT using SimMIM algorithm[66] on the all collected 96K CTs, and then conduct further supervised fine-tuning on the 6K CTs with 150K labeled segmentation masks. The image encoder, denoted as $\mathcal{F}_{\text{IE}}(*,\bm{\theta}_{\text{IE}})$, takes a volumetric image $\bm{x}\in\mathcal{R}^{C\times D\times H\times W}$ as input. Firstly, it splits $\bm{x}$ into a set of patches, denoted as $\bm{x}_{\text{patch}}\in\mathcal{R}^{N\times(C\times P_{D}\times P_{H}\times P_{W})}$, where $N=\frac{D\times H\times W}{P_{D}\times P_{H}\times P_{W}}$. $P_{D}$, $P_{H}$ and $P_{W}$ are the size of patch. These patches are then fed into the network, which outputs an embedding $\bm{z}_{\text{image}}=\mathcal{F}_{\text{IE}}(\bm{x}_{\text{patch}},\bm{\theta}_{\text{IE}})$, $\bm{z}_{\text{image}}\in\mathcal{R}^{N\times F}$. $F$ represents the feature dimension, which is set to 768 by default in this paper.
One main limitation of traditional segmentation models is that the models learn dataset-specific labels encoded as integers which cannot generalized to new datasets or tasks, limiting their real-world applications. We enable universal segmentation across datasets by leveraging the text prompt. We employ the text encoder from CLIP model[45] to encode the input text prompt, as CLIP[45] has been trained to align image and text on web-scale image-text pairs. We denote the text prompt encoder as $\mathcal{F}_{\text{TE}}(*,\bm{\theta}_{\text{TE}})$. Given a word or phrase as prompt, we complete it using the template $\bm{s}=$‘A computerized tomography of a [text prompt]’[35]. $\bm{s}$ is then tokenized into $\bm{t}$. The text encoder accepts $\bm{t}$ as input and outputs the text embedding $\bm{z}_{\text{text}}=\mathcal{F}_{\text{TE}}(\bm{t},\bm{\theta}_{\text{TE}})$, where $\bm{z}_{\text{text}}\in\mathcal{R}^{F}$. We freeze the off-the-shelf text encoder during training since the text data in CT datasets is a small amount.
Following SAM[31], we use the positional encoding[60] for point prompt $\bm{p}$ and bbox prompt $\bm{b}$ and obtain the point embedding $\bm{z}_{\text{point}}\in\mathcal{R}^{F}$ and bbox embedding $\bm{z}_{\text{bbox}}\in\mathcal{R}^{F}$. We concatenate the embeddings of three kinds of prompts as $\bm{z}_{\text{prompt}}=\mathcal{F}_{\text{PE}}(\bm{p},\bm{b},\bm{s},\bm{\theta}_{\text{PE}})=[\bm{z}_{\text{point}},\bm{z}_{\text{bbox}},\bm{z}_{\text{text}}]$.
After obtaining the image embedding $\bm{z}_{\text{image}}$, prompt embedding $\bm{z}_{\text{prompt}}$ and text embedding $\bm{z}_{\text{text}}$, we input them to the mask decoder and predict the mask $\bm{p}=\mathcal{F}_{\text{MD}}(\bm{z}_{\text{image}},\bm{z}_{\text{prompt}},\bm{z}_{\text{text}},\bm{\theta}_{\text{MD}})$. We use self-attention and cross-attention[63] in two directions to blend the image embedding and prompt embedding, and then employ the transposed convolutions and interpolation operations to generate masks. Since the text embedding is the key to universal segmentation and it is also harder to learn the correlation between text and volumetric regions, we reinforce the text information by introducing a parallel text input $\bm{z}_{\text{text}}$ beside the joint prompt embedding $\bm{z}_{\text{prompt}}$. We further compute a similarity matrix between the up-scaled embedding from the transposed convolution output and the text embedding in the mask decoder. The element-wise multiplication of the similarity matrix with the mask prediction is applied before interpolation, after which the model outputs the masks.
4.3 Prompt generation
SegVol can accept multiple prompt types, including individual point prompts, bbox prompts, and text prompts, and also their combinations. To make full use of the segmentation training data, we generate kinds of prompts for each datum. Then, the prompt and mask pairs are used to compute the training loss. SegVol supports ‘point’ prompts, ‘bbox’ prompts, and ‘text’ prompts.
The point prompt is built from ground truth or pseudo masks, consisting of three kinds of points, namely, positive point, negative point, and ignore point. Positive point means that it is within the target mask region, while negative points are those outside. The ignore points are used for input completion, which will be disregarded by the model so that the point prompt has the same length.
The bbox and point prompts can be generated by sampling points based on the ground-truth segmentation masks, while text prompts are constructed based on their category names. As pseudo masks produced by the unsupervised FH algorithm[11] do not have the semantic information, we only use point and bbox prompts when training with pseudo masks.
4.4 Loss function
We apply SimMIM algorithm[66] to pre-train the image encoder of SegVol using the masked image modeling loss $\mathcal{L}_{\text{pre-training}}(\bm{\theta}_{\text{IE}};\mathcal{D}_{1})$. The loss function is as follows:
$\mathcal{L}_{\text{pre-training}}(\bm{\theta}_{\text{IE}};\mathcal{D}_{1})=\frac{1}{\Omega(\bm{x}_{\text{M}})}||\bm{y}_{\text{M}}-\bm{x}_{\text{M}}||_{1},$ (1)
where $\bm{x},\bm{y}\in\mathbb{R}^{D\times H\times W}$ are the input voxel values and predicted values, respectively. M denotes the set of masked voxels, $\Omega(\cdot)$ is the number of elements, and $\mathcal{D}_{1}$ is the pre-training dataset.
We combine the binary cross-entropy (BCE) loss and Dice loss as the supervised fine-tuning loss function $\mathcal{L}_{\text{fine-tuning}}(\bm{\theta};\mathcal{D}_{2})$ to train the model with parameters $\bm{\theta}$, where $\bm{\theta}=[\bm{\theta}_{\text{IE}},\bm{\theta}_{\text{PE}},\bm{\theta}_{\text{MD}}]$ and $\mathcal{D}_{2}$ is the supervised fine-tuning dataset. The loss function is as follows:
$\mathcal{L}_{\text{BCE}}(\bm{\theta};\mathcal{D}_{2})=$ $-\mathbb{E}_{(\bm{x},\bm{y})\sim\mathcal{D}_{2}}[\langle\bm{y},\log(\mathcal{F}(\bm{x},\bm{\theta}))\rangle+$ (2)
$\langle 1-\bm{y},\log(1-\mathcal{F}(\bm{x},\bm{\theta}))\rangle]$
$\mathcal{L}_{\text{Dice}}(\bm{\theta};\mathcal{D}_{2})=1-\mathbb{E}_{(\bm{x},\bm{y})\sim\mathcal{D}_{2}}[\frac{2\langle\bm{y},\mathcal{F}(\bm{x},\bm{\theta})\rangle}{\|\bm{y}\|_{1}+\|\mathcal{F}(\bm{x},\bm{\theta})\|_{1}}]$ (3)
The detailed supervised fine-tuning training loop of SegVol is presented in Supplementary Algorithm 1.
4.5 Zoom-out-zoom-in mechanism
Compared to 2D slides, volumetric data has a remarkably large number of voxels meanwhile small segmentation targets relatively. Naively down-sampling the original data will cause serious information loss and thus inferior performance. Diving the large volumetric data into small cubes and conquering each separately is computationally expensive and also suffers from information loss. To reduce the computational cost while preserving the details of the Region of Interest (ROI), we design the zoom-out-zoom-in mechanism consisting of multi-size training and zoom-out-zoom-in inference.
To adapt various sizes of volumetric data and enable the zoom-out-zoom-in inference, we construct two kinds of training data. One is to resize the large-size CT to adapt the model’s input size and obtain the training data of the zoom-out view. The other one is to crop the original large-size CT into cubes with the model’s input size. In this way, we obtain the training data of zoom-in view.
During the zoom-out-zoom-in inference, we first zoom out and implement global inference. Given a large volumetric image, it is resized and then fed into the SegVol model. After obtaining the global predicted segmentation mask based on the user’s prompt, we locate the region of interest (ROI) and zoom in, namely, crop it from the original-size image. We apply a sliding window on the cropped region and implement more precise local inference. We adapt the input prompt for the local inference, since the original point and bbox prompts input by the user may not be applicable in the local inference region when zoom-in. Specifically, we ignore the positive or negative points of the local region. Similar to the training bbox prompt generation in Sec. 4.3, we generate the local bbox prompt by considering the global predicted mask in the local region as the pseudo mask. Finally, we fill the ROI region of the global segmentation mask with the local segmentation mask. The zoom-out-zoom-in mechanism realizes both efficient and precise inference simultaneously. The detailed procedure of zoom-out-zoom-in mechanism is provided in Supplementary Fig. 1 c and d.
4.6 Evaluation metrics
During the training and internal validation process, each subset of the joint dataset is divided into 80% training data and 20% internal validation data. To ensure the absence of any data leaks, hash value is utilized to compare between the validation set and the training set. And during the external validation process, the model’s parameters are all frozen.
5 Data and code availability
The training, internal, and external validation datasets used in this study are publicly accessible and licensed. Please refer to their original papers for the details of these datasets. The download links are provided in Supplementary Table 2.
Our training code, inference code, and model weights have been publicly available in https://github.com/BAAI-DCAI/SegVol. The online running demonstration is provided in https://huggingface.co/spaces/BAAI/SegVol.
Acknowledgement
Funding: This work is funded by the National Key R&D Program of China (2021ZD0111102) and NSFC-62306046.
Supplementary Materials

### v5

@@ 3.1 Experimental Setup
Figure 2: Violin plots for quantitative comparison experiment results of SegVol and SAM-like interactive methodskirillov2023segment ; cheng2023sam ; wang2023sammed3d ; ma2023segment . The vertical axis represents the Dice score.
3.2 Compared with SAM-like Interactive Methods
@@
Several efforts have been made to construct a SAM-like interactive medical image segmentation model. However, some of these works, such as MedSAMma2023segment and SAM-MED2Dcheng2023sam , focus on 2D tasks and cannot process 3D input directly. The other 3D-based methods, such as SAM-MED3Dwang2023sammed3d , only support small cropped input and do not support semantic-prompt segmentation, which are still far from building a comprehensive foundation model for volumetric medical image analysis.
Competitors and configures.
In this experiment, MedSAMma2023segment and SAM(bounding box)kirillov2023segment use bounding box prompts. SAM(5 clicks)kirillov2023segment , SAM-MED2Dcheng2023sam and SAM-MED3Dwang2023sammed3d use point prompts and a five-step correction procedure, which means that the point prompt in each step will be given according to the previous-step output and ground truth, rather than giving all at once. In this experiment, SegVol uses bounding box and text prompt which performs better than other kinds of prompt combinations. Detailed ablation study on prompt combination is demonstrated in Figure 3 (b). In addition, we compare SegVol with traditional task-specific segmentation models, e.g., 3DUX-NETlee20233d , SwinUNETRhatamizadeh2022swin , and nnU-Netisensee2021nnu , in Section C, though the direct comparison is unsuitable due to the different settings and objectives.
Testing data.
To compare with these SAM-like interactive segmentation models, we evaluate the models on 1,778 cases from the validation set of AMOS22ji2022amos , the whole novel annotated set of Universal Lesion Segmentation Challenge 23(ULS23)uls23 , and the released labeled set of SegTHORlambert2019segthor . The validation set of AMOS22 contains 120 cases annotated with 15 major organs. The novel annotated ULS23 dataset is composed of three subsets, namely, DeepLesion3D, Radboudumc Bone, and Radboudumc Pancreas. The DeepLesion3D subset contains 200 abdominal lesions, 100 bone lesions, 50 kidney lesions, 50 liver lesions, 100 lung lesions, 100 mediastinal lesions, and 150 assorted lesions cases. There are 744 bone lesion cases in the Radboudumc Bone subset and 124 pancreas lesion cases in the Radboudumc Pancreas subset. The 40 cases from SegTHOR, which are contoured manually by an experienced radiotherapist, focus on the heart, trachea, aorta, and esophagus that surround the tumor and must be preserved from irradiations during radiotherapy.
Quantitative results.
The quantitative results of comparative experiments are shown in Table 2, which verify our method is the best in most of the tasks including both lesions and organs, compared to other SAM-like interactive modelskirillov2023segment ; cheng2023sam ; wang2023sammed3d ; ma2023segment . Specifically, our method outperforms the second-ranked SAM-MED2D on the AMOS22 dataset by a significant improvement of 19.25% (average Dice score). On the SegTHOR dataset, our method surpasses the runner-up – SAM-MED3D by an average Dice score improvement of 4.28%. The ULS23 dataset, characterized by small patch-like masks, presents a unique challenge. In this scenario, SegVol still exhibits good performance, comparable to MedSAM, which excels in using bbox prompts for segmenting small objects. We visualize the Dice score distributions of all methods in all the tasks as violin plots, depicted in Figure 2. More detailed results and visualization are present in Section C.
Table 3: Ablation experiment on the zoom-out-zoom-in mechanism.
Mechanism
Dice Score Avg. $\uparrow$
Time Per Case Avg. $\downarrow$
Resize 0.4509 65 ms
Sliding window 0.6529 3331 ms
Zoom-out-zoom-in 0.7298 190 ms
Figure 3: (a) The performance of SegVol improves as the training data scales up. (b) The quantitative experimental results on 19 anatomical segmentation tasks of split 20% test data demonstrate that using the combination of semantic and spatial prompts can achieve better performances.
3.3 Ablation Studies
Zoom-out-zoom-in mechanism.
One of the key designs of SegVol is the zoom-out-zoom-in mechanism. We compare it with the intuitive resize strategy and the popular sliding window algorithm on the split 20% test data, 48 cases covering 15 major organs with a variety of sizes, belonging to the AMOS22ji2022amos dataset. Two evaluation dimensions, i.e., performance (Dice score) and inference time cost (per case), are compared, as shown in Table 3. The zoom-out-zoom-in mechanism achieves the best average Dice score and a very competitive inference speed compared to the simple resize strategy. The reason for computational cost reduction is that the traditional sliding window method requires scanning the entire 3D CT and processing thousands of windows. In contrast, the proposed zoom-out-zoom-in mechanism only requires one global inference of 3D CT and then scanning the ROI with dozens of windows. Detailed experiment results are shown in Section C.
Scaling up training data.
@@ 3.3 Ablation Studies
The success of scaling up training data has been witnessed in multiple computer vision tasks kirillov2023segment ; radford2021learning . We conduct an ablation study to investigate the importance of scaling up training images and masks. The split 20% test data of BTCV datasetlandman2015miccai , which includes 13 main organs, is set as an anchor to evaluate the model trained separately on 1, 2, and 8 datasets for 500 epochs, as well as the final model trained on 25 datasets. The detailed results are shown in Figure 3 (a). As a lightweight model, the performance of SegVol is weak when only one dataset is used. However, with the increase of training data, the Dice score increases rapidly, especially in the text prompt setting. The results indicate that our method is scalable and better performance can be achieved if more training data is available.
Multi-prompt combination.
As a universal model, our approach achieves precise segmentation for over 200 organs, tissues, and lesions using both spatial and semantic prompts. In Figure 3 (b), we quantitatively analyze the mutually supportive relationship between semantic-prompt and spatial-prompt in 19 segmentation tasks of the 20% split test data. On the one hand, spatial-prompt allows the model to locate the specific part in the 3D space. According to Figure 3 (b), the average Dice score of the ‘bbox+text’ prompt is boosted by 5.85% compared to the ‘text’ prompt on average. On the other hand, semantic-prompt clarifies the reference to the anatomical structure, eliminating the ambiguity of spatial-prompt and the plausible masks of multiple categories. This is reflected in Figure 3 (b) as the average Dice score of ‘point+text’ prompts is 4.62% higher than using ‘point’ prompts alone. Spatial and semantic prompts mutually support each other, ultimately endowing the model with powerful segmentation capabilities.
3.4 Case Studies
Disambiguation via semantic-prompt.
It is a notorious problem in interactive segmentation that one spatial-prompt may correspond to multiple plausible outputs kirillov2023segment . As illustrated in the images on the top left in Figure 4, three of them correspond to three anatomical concepts, namely, kidney tumor, left kidney, and the whole kidneys, while they are all plausible to the same point prompt. Similarly, in the bottom left three images, the bounding box selects the region of the liver. However, liver tumors, hepatic vessels, and the liver itself are also plausible target structures. In these cases, SAM chooses to return multiple masks to match different levels of plausible results. Unlike SAM’s solution, we use semantic-prompt to clarify the targets. As shown in Figure 4, the captions below the images are the text prompts, and the masks in the images are the predictions of SegVol, which show that semantic-prompt can effectively disambiguate the spatial-prompt.
Identifying the spatial-prompt segmentation.
Furthermore, we study the capability of SegVol to identify the semantic category of the spatial-prompt results. Figure 5 reveals that SegVol can give accurate semantic categories based on the spatial-prompt results. In the top left image in Figure 5, the spatial-prompt on the liver results in a 0.997 prediction score for the liver. The top right image in the sub-figure shows if the spatial-prompt is the point on the liver tumor, SegVol will output a 0.619 prediction score for the tumor category and a 0.339 prediction score for the liver based on the spatial relationship of liver tumor and liver. We implement this identification experiment by decoding the semantic prompts from a category set. The softmax function is applied to the decoding results to get the prediction probabilities of different categories. The probabilities on the initial predicted mask, driven by the spatial-prompt, are used to calculate the final classification result.
Figure 4: The four cases demonstrate that semantic-prompt can clarify the ambiguity of spatial-prompt and avoid multi-plausible outputs. Each image shows the segmentation result of SegVol using the spatial-prompt, i.e. point or bounding box, and semantic-prompt, i.e. the caption below the image.
Figure 5: We identify the semantic categories of the spatial-prompt segmentation results. Each image shows the spatial-prompt and the mask prediction. The bar charts rank the top 8 semantic categories with the highest classification probabilities. The results show that SegVol is capable of identifying the anatomical category of the segmentation mask using spatial prompts.
@@ 4 Discussion
Scalability.
The scaling law of foundation models has been verified in multiple CV and NLP tasks. Since SegVol uses a transformer-based architecture and self-supervised pre-training algorithm, it has strong data and architecture scalability. In this work, we achieve the success of scaling law in 3D medical segmentation by the design of universal prompts and pseudo masks for joint learning on datasets with inconsistent annotations. The ablation study of scaling up training data shows that 1) the performance improves significantly with more training data in the 3D segmentation task, 2) SegVol has not yet reached its ceiling if more training data is provided. We believe the performance of SegVol can be continuously improved when more data and computational resources are used.
Generalizability to unseen modality.
Although we develop SegVol on Computed Tomography (CT) data due to its advantages of easy acquisition, wide usage, and high resolution, we find that SegVol can generalize to other medical image modality, like MRI. Namely, the SegVol model trained only on CT data can be used to segment MRI with semantic and spatial prompts. This emerging ability demonstrates that our foundation model understands the anatomical structure of human body. We provide detailed experiments and analysis of this generalizability in Section C. The impressive generalizability makes SegVol a versatile tool in medical image analysis. We leave the joint training of SegVol on multi-modality data as the future work.
Limitations.
Although SegVol shows remarkable semantic-prompt segmentation performance, there still remains gap between it and the referring volumetric segmentation. A promising solution is to construct the referring segmentation data with diverse semantic and spatial prompts, and then train SegVol on it. We leave it as the future work. More discussions can be found in Section E.
Broader impact.
We contribute a foundation model for universal and interactive volumetric medical image segmentation, which can benefit numerous clinical study and applications. As a foundational research work, we do not see any obvious negative societal impact of the proposed method and model.
5 Conclusion
In this paper, we propose SegVol, a universal and interactive volumetric medical image segmentation model, supporting both spatial-prompt and semantic-prompt segmentation of more than 200 anatomical categories. We construct a large-scale dataset, which consists of 90K unlabeled CTs and 25 open-source medical datasets, to train the foundation model. We design the zoom-out-zoom-in mechanism to facilitate efficient and precise inference in the region of interest. Extensive experiments on 22 segmentation tasks demonstrate the outstanding performance of our method. Detailed ablation studies are also carried out to prove the effectiveness of the zoom-out-zoom-in mechanism, dataset scale, and multi-prompt combination strategy. As a foundation model, we believe that SegVol will advance the volumetric medical segmentation and benefit numerous downstream tasks.
Acknowledgements.
This work is funded by NSFC-62306046.
Appendix A Dataset Details and Availability
In this work, we collect 25 open-source datasets for supervised fine-tuning SegVol and some external open-source datasets specifically for comparative experiments. The detailed information on anatomical categories and dataset scales of these open-source datasets is shown in Table 4. The availability of these datasets is demonstrated in Table 5. Additionally, to avoid privacy concerns, we collect 90K unlabeled CTs from publicly accessible professional medical websites: https://radiopaedia.org/.
The collected segmentation datasets include major regions of the human body, i.e., the head, neck, thorax, abdomen, and pelvis, comprising over 200 categories of organs and tissues, and 28 lesion tasks from different benchmarks. The detailed categories information are shown in Figure 6 and some representative samples are shown in Figure 7.